Which atom has the smallest number of neutrons? carbon-14 nitrogen-14 oxygen-16 fluorine-19 neon-20 Please explain why the answer is correct :)
Answer #1

Ans : nitrogen -14

The number of neutrons = mass number - atomic number

for carbon : 14 - 6 = 8

for nitrogen : 14 - 7 = 7

for oxygen = 16 - 8 = 8

for fluorine = 19 - 9 = 10

for neon = 20 - 10 = 10

the smallest number of neutrons are present in nitrogen-14 and hence the answer.